OpenStudy (anonymous):

The amount of a radioactive material changes with time. The table below shows the amount of radioactive material f(t) left after time t. t(hours) 0 1 2 f(t) 100 50 25 Which exponential function best represents the relationship between f(t) and t? f(t) = 0.5(100)t f(t) = 100 (0.25)t f(t) = 0.25(50)t f(t) = 100(0.5)t

OpenStudy (anonymous):

@pratyush5 I hate to bother you again but you seem like the only person that can answer my math qs

OpenStudy (pratyush5):

Sorry to disappoint but i havent studied functions yet. :(

OpenStudy (pratyush5):

@amoodarya maybe he can help

OpenStudy (anonymous):

ok but thank you for trying. :D

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

@amoodarya could you plz help me out? :)

OpenStudy (amoodarya):

t(hours) 0 1 2 f(t) 100 50 25 do you mean ? 10 for 0 50 for 1 25 for 2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

100 for 0

OpenStudy (amoodarya):

suppose f(t)= a * (b)^c now put point in it to find a ,b, c

OpenStudy (anonymous):

What does a and b represent?

OpenStudy (amoodarya):

it easier to say f(t)= a*(1/4)^b because ratio is 1/4 =0.25

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

f(t) = 100 (0.25)t

OpenStudy (anonymous):

so would f(t) = 100 (0.25)t be my answer? :)

OpenStudy (amoodarya):

put (0,100) 100=a* (1/4)^0 100=a*1 a=100 then put (1,50) 50 =a*(1/4)^(b) 50=100 *(1/4)^b 1/2=(1/4)^b

OpenStudy (amoodarya):

so b=1/2 and f(t)=100 (1/4)^(t/2) it means f(t)=100(0.5)^t

OpenStudy (anonymous):

ok i think i understand it now, thank you very much! :D

OpenStudy (amoodarya):

you did not type it correct but it 4th choice

OpenStudy (anonymous):

ok thank you so much :D